OpenStudy (anonymous):

Keiko is saving money to buy a game. The game costs $9 , and so far she has saved one-third of this cost. How much money has Keiko saved?

OpenStudy (kamibug):

$3

OpenStudy (anonymous):

can u write an eqation on solving it plz

OpenStudy (kamibug):

9/1 * 1/3 = 9/3 = 3

OpenStudy (anonymous):

ok thank u so much

OpenStudy (kamibug):

no problem =)